OpenStudy (anonymous):

9 (less than or equal to sign)negative9+d I think the answer is d (less than or equal to sign) -18

OpenStudy (anonymous):



OpenStudy (loser66):

\[9\leq -9+d\] +9 both sides, you get d, no change for the sign, it is still \(\leq\)

OpenStudy (anonymous):

Wouldn't it be minus 9 on bot sides instead of +?

OpenStudy (anonymous):

both not bot @Loser66

OpenStudy (loser66):

no, + , not - because : when +9 to the right hand side, that +9 and -9 cancel out to get d

OpenStudy (anonymous):

so the answer is just d?

OpenStudy (loser66):

yes, you must have d .... something ,

OpenStudy (anonymous):

18

OpenStudy (loser66):

oh boy!! I am ambiguous.!! OK, \[9\leq -9+d\\9+9 \leq -9 +d+9\\18\leq d\]

OpenStudy (anonymous):

thank,mind check a few more for me?

OpenStudy (anonymous):

i'll tag you in it

OpenStudy (loser66):

Let see if it is not above my head, I will try